Assume that U is a plane.

Find out whether or not the following vectors lie in U:

(10.1) ~u =< 3.8, 1 >, ~v =< −4, 1, 1 > and w~ = −~v

(10.2) ~u =< 3.8, 1 >, ~v =< −4, 1, 1 > and w~ = ~u − ~


1
Expert's answer
2021-06-09T08:47:17-0400

Vectors lie in U if

"u\cdot(v\times w)=0"


1.

"v\times w=\begin{vmatrix}
 i & j&k \\
 -4 & 1&1\\
4&-1&-1
\end{vmatrix}=0"

All vectors are in the plane U.


2.

"v\times w=\begin{vmatrix}
 i & j&k \\
 -4 & 1&1\\
-3&-8&-1
\end{vmatrix}=7i-7j+35k"

"u\cdot(v\times w)=(3,8,1)\cdot(7,-7,35)=21-56+35=0"

All vectors are in the plane U.